OpenStudy (anonymous):

Can anyone help with a Induction heating problem. I am basically from Mechanical background, working over a Industrial problem of Induction heating . We have a coil of 54 number of turns, with voltage ( 650V) and Power ( 150 KW) applied across its Junction. Can anyone help me to calculate the voltage and power available for single turn of the coil. Please note that the Coil is consist of 4 different radius ( decreasing from starts to end) with the following number of turns 18 17 12 7.

OpenStudy (mos1635):

the 4 diferent radius is needed

OpenStudy (anonymous):

Radius are as follow: 78 73 68 63

OpenStudy (anonymous):

In mm

OpenStudy (experimentx):

do you have a figure??

OpenStudy (anonymous):

Not now, But I will send it to you tomorrow. Meanwhile, you can assume the four coil connected to each other forming a long solenoid of 900mm.

OpenStudy (anonymous):

See If you can come out with suggestions.

OpenStudy (experimentx):

I think we need radius of the turns!!

OpenStudy (anonymous):

Radius are as 78 mm( 7 turns) 73 mm(18 turns) 68 mm(17 turns) 63mm( 12 turns)

OpenStudy (anonymous):

for total of 54 turns

OpenStudy (anonymous):

Above are the diameters

OpenStudy (experimentx):

find the total length of all turns ..

OpenStudy (anonymous):

total lenght is 900 mm

OpenStudy (anonymous):

Pl tell : 1. Is inductance of the turns has effect over the voltage ?? 2. If the cross section of the coil is rectangular and I have given your the mean dia i.e. \[Dmean=( Doutside + Dinside)/2\] Then again we have to follow the same procedure. 3. And what about the power distribution ??

OpenStudy (mos1635):

78 mm( 7 turns) 73 mm(18 turns) 68 mm(17 turns) 63mm( 12 turns) are these radius OR diametres ??

OpenStudy (experimentx):

Sorry ... I think i misunderstood the question.

OpenStudy (vincent-lyon.fr):

In order to know what the induced emf is, we need to know the frequency of the generator.

OpenStudy (vincent-lyon.fr):

What is supposed to be heated? The coil itself or something that you would put inside it?

OpenStudy (anonymous):

A cylindrical wire(12.5 mm diameter) of infinite length is running at a finite speed ( 1.5m/s) inside the coil. The frequency is 5 Khz. Diameters are given above not radius.

OpenStudy (anonymous):

i need a figure plz\

OpenStudy (anonymous):

cant we just separate all the four kind of coils as being in series and evaluate voltage available, through first calculating for each kind and then taking a mean (by considering it as a frequency distribution data)?? if the condition imposed is that flux through each coil is same, then we can forget the difference in radius and the associated frequency distribution and just calculate total length calculation (which is a trivial calculation) and using total resistance,and thus find average power and voltage for each coil...

OpenStudy (anonymous):

Here is the solution : Please give your opinion Inductance of the whole coil would be = Xl1+Xl2+Xl3+Xl4 where Xl1= 2*pi*f*L1 L1= U0*N*N*A/l1 Total impedance (Z)= Xl ( Ignoring R^2 << Xl ^2) Volatage across N1 turns = V/Z *( Xl1)

OpenStudy (vincent-lyon.fr):

@Ishant: You said : "Inductance of the whole coil would be = Xl1+Xl2+Xl3+Xl4 where Xl1= 2*pi*f*L1 L1= U0*N*N*A/l1" this is certainly a good approximation. First two lines are ok The expression you give for L1 is supposed to be valid for an long (infinite) coil, which is not the case here at all, but value must be not too far from that.

OpenStudy (anonymous):

@vincent: I have taken help from here : http://hyperphysics.phy-astr.gsu.edu/hbase/hframe.html Please see and let me know, If this is Ok.